Grade 12General Physics

​​A particle moves in a straight line.its speed(v)increases lenearly with time. If the initial speed of the particle is v and its speed at t=4s is 2v then ∫04vdt is equal to

Santhoshi
7 Years agoGrade 12

1 Answer

Khimraj
7 Years ago
speed V = v  + ct
where c is consatnt
so at t= 4
2v = v + c*4
so c = v/4
so V = v + vt/4
then \int_{0}^{4}Vdt = \int_{0}^{4}(v + vt/4)dt = [vt + vt^{2}/8]_{0}^{4} = 4v + 2v = 6v
Hope it clears.......................